Length of DNA molecule 1700A. If adenine = 20. The number of 2Hbonds are
A)  60   B)  30  C)  100  D)  130

Grade:12

2 Answers

Arun
3 years ago
Length of a base pair is 3.4 Å. Hence a DNA segment of length 170 Å has 170/3.4 = 50 base pairs.
It is given in question that 20% of DNA is cytosine. Cytosine forms triple hydrogen bond with guanine resulting in a base pair Hence, an equal percentage ( 20% ) of guanine is present too. Based on this it concluded that 40% of base pairs ( 50×40% = 20 base pairs ) are cytosine-guanine. Each of them has triple bond hence 20×3 = 60 hydrogen bonds.
While rest, 60% ( 50×60% = 30 base pairs ) are adenine-thymine. Each of them has a double bond hence 30×2 = 60 hydrogen bonds. 
 
Hence this dsDNA segment has 60+60 = 120 hydrogen bonds.
Vikas TU
3 years ago
In three turns there will be 30 base pairs of DNA or 60 nucleotides. So, 12 nucleotides are thymine. thymine is double bonded with hydrogen bonds with adenine so 12 nucleotides are adenine according to Chargaff's rule. So it will have 24 hydrogen bonds